OpenStudy (anonymous):

f(x)=-2x^2+4x+3, use completing the square method to put the function into f(x)=a(x-h)^2+k

OpenStudy (anonymous):

=-2(x²-2x)+3 =-2(x²-2x+1-1)+3 =-2(x²-2x+1)+2+3 =-2(x-1)²+5

OpenStudy (anonymous):

thanks